Solve the equations for x.


25x squared = 13



2(x-5)squared = 3

Answer by checkley71(8403)

25X^2=13
X^2=13/25
X=SQRT13/25
X=.72 ANSWER.
--------------------------------------
2(X-5)^2=3
(X-5)^2=3/2
X-5=SQRT3/2
X=5+1.2247
X=1.2247-5
X=-3.7753 ANSWER.